### v1

@@ 6 Experiments
In this section, we evaluate the effectiveness of flopping-equivariant networks. We first briefly discuss the setting of the experiments. Then we compare the efficiency and accuracy of equivariant versions of ResMLPs, ViTs and ConvNeXts from Section 4 to their non-equivariant counterparts. For a given architecture $X$, the equivariant version is $\mathcal{E}(X)$. $\mathcal{E}(X)$ always has around half the number of trainable parameters and FLOPs of $X$ due to the block-diagonalization of the weight matrices (2). We will release code and weights on GitHub.
Dataset. We benchmark our model implementations on the ImageNet-1K dataset, which includes 1.2M images evenly spread over 1,000 object categories.
Hyperparameters. We use the same training recipes as the baselines. The complete set of hyperparameters can be found in Table 2.
Implementation. The major building blocks of modern deep learning, including linear layers, convolution layers and attention layers, are all implemented using general matrix multiplications (GEMMs). Modern GPUs enable efficient GEMMs, typically implemented in a tiled fashion, processing blocks of the output matrix in parallel. Consequently, reducing the linear layers into block diagonals does not hinder GPU utilization, as the same tile-GEMMs can be used for the full product (1) or the block-diagonal (2) (with the latter requiring half as many). We use the PyTorch Paszke et al. (2019) compiler to obtain efficiently running networks, but writing custom GPU-code would likely further improve the throughput.
@@ 6.1 Results and discussion
In prior work, image classification is often contextualized as a trade-off between accuracy and measures of model complexity, such as FLOPs and number of parameters. To that end, in Table 1, we highlight both the theoretical complexity, i.e. model size, and FLOPs, as well as the empirical speed as measured by the throughput (images processed per second), and the Top-1 classification accuracy on ImageNet-1K. Our results demonstrate that as we increase the scale of the networks, flopping-equivariant networks achieve competitive classification accuracy while requiring half the FLOPs, as highlighted in Figure 1, and showcase higher throughput compared to baseline implementations. The hybrid model $(\text{ViT-H})$-H ) achieves the best accuracy of all models, aligning with the results for hybrid models by Weiler & Cesa (2019). One reason the smaller equivariant models underperform in terms of accuracy is that they are quite limited in the number of parameters—the regularization methods in the training recipes for the corresponding non-equivariant models (with more parameters) may be too heavy. We leave the compute-heavy task of optimizing the training recipes for future work, as we believe that our experiments clearly demonstrate the main point of this paper—that equivariant networks are scalable.
Table 1: Classification with Imagenet1k training. We contrast our flopping-equivariant networks to the originals using the same training recipes. The equivariant version of architecture X is denoted $\mathcal{E}(\text{X})$. $(\text{ViT-H})$-H ) is a hybrid model with equivariant layers for the first half of the network. The baselines are not rerun, instead we show the results reported in their respective papers. The throughput and peak memory are measured on a single A100-40GB GPU with batch size fixed to 64 and compiled networks running mixed precision forward passes with no gradients.
@@
MLP-Mixers (ResMLP Touvron et al. (2023))
@@
Vision Transformers (DeiT III Touvron et al. (2022))
@@
Convolutional Networks (ConvNeXt Liu et al. (2022))
@@
As shown in Figure 6, the relative throughput improvement of the equivariant networks becomes more pronounced as the size of the model increases, which can be attributed to a higher proportion of time spent on the linear layers, with the computational overhead from less optimized kernels becoming less significant. This is best illustrated by the superior speed of the equivariant ResMLP models for which linear layers dominate the computation. Our networks have no throughput improvement for attention or depthwise convolutions, which, despite not contributing the most FLOPs, are sometimes comparable in runtime to the dense layers. Future more efficient implementations of these layers, and further scaling of the embedding dimension, would pronounce the gain in throughput for the equivariant models.444 The reader will notice a flattening out of the rate of improvement between ViT-L and -H. ViT-H uses patch size $14\times 14$ while the other ViTs use patch size $16\times 16$. Smaller patch sizes means more patches and more compute in the attention layers.
6.2 Experimental limitations
[…]
@@ 7 Conclusion
The considered task of upright image classification is in a sense the least interesting for equivariant networks as the symmetry group is small and the output is invariant. We believe that future efforts should be directed towards equivariant networks as general visual feature extractors. Classic and recent works have already demonstrated the utility of having equivariant features for downstream vision tasks Lowe (2004); Loy & Eklundh (2006); Lee et al. (2023); Bökman et al. (2024); Garrido et al. (2024).
@@ Appendix A Groups, representations and Schur’s lemma
This section is not required reading to understand the main takeaway of the paper. It serves to describe how the content of Section 3 can be formalized and generalized using group theory. The theory content of this section is common knowledge in geometric deep learning Bronstein et al. (2021); Gerken et al. (2023); Weiler et al. (2023) and part of standard representation theory courses Serre (1977). Our contribution is to highlight the computational aspect more than has perhaps been done in previous work.
@@
Following the geometric deep learning blueprint Bronstein et al. (2021), to design flopping-invariant neural networks $f=f_{N}\circ\cdots\circ f_{2}\circ f_{1}$, we enforce each network layer $f_{i}$ to be flopping-equivariant. The last layer is designed be flopping invariant, which is the special case of (11) where $\rho_{2}(g)\equiv I$. Designing an invariant neural network hence involves a choice of intermediate group representation $\rho_{i}$ for the output of each layer. We can write down all possible choices of $\rho_{i}$ in terms of fundamental building blocks called irreducible representations.
@@
Let $\varrho_{1}$ and $\varrho_{2}$ be two real irreps of the finite group $G$.
@@
The dimension in case 2 depends on whether the irrep $\varrho_{1}$ is of so-called real, complex or quaternary type Boardman (2007).
@@
DeiT III Touvron et al. (2022) ResMLP Touvron et al. (2023) ConvNeXt Liu et al. (2022)
@@ 3 Augment
MLP-Mixers (ResMLP Touvron et al. (2023))
@@
Vision Transformers (DeiT III Touvron et al. (2022))
@@
Convolutional Networks (ConvNeXt Liu et al. (2022))
@@ Appendix B Experimental setting
Software versioning. Our experiments build upon PyTorch Paszke et al. (2019) and the timm Wightman (2019) library. We enable mixed-precision training using the deprecated NVIDIA library Apex, this is to mirror the training recipes of the benchmarks as closely as possible. To enable PyTorch’s compiler, we use a modern version ($\geq 2.0$). Specifically, we use PyTorch 2.5.1 with CUDA 11.8.
@@
    Flash Attention Dao (2023), used by the ViTs.
@@ Appendix C ResMLP linear layers over patches
This calculation is an example of a Clebsch-Gordan product or tensor product of representations, which has been used in some prior designs of equivariant networks Kondor et al. (2018).

### v2

@@ 6 Experiments
In this section, we evaluate the effectiveness of flopping-equivariant networks. We first briefly discuss the setting of the experiments. Then we compare the efficiency and accuracy of equivariant versions of ResMLPs, ViTs and ConvNeXts from Section 4 to their non-equivariant counterparts. For a given architecture X, the equivariant version is $\mathcal{E}(\text{X})$. $\mathcal{E}(\text{X})$ always has around half the number of trainable parameters and FLOPs of X due to the block-diagonalization of the weight matrices (2). We will release code and weights at github.com/georg-bn/flopping-for-flops.
Dataset. We benchmark our model implementations on the ImageNet-1K dataset (Deng et al., 2009; Russakovsky et al., 2015; Recht et al., 2019), which includes 1.2M images evenly spread over 1,000 object categories.
Hyperparameters. We use the same training recipes as the baselines. The complete set of hyperparameters can be found in Table 2 in the appendix.
Implementation. The major building blocks of modern deep learning, including linear layers, convolution layers and attention layers, are all implemented using general matrix multiplications (GEMMs). Modern GPUs enable efficient GEMMs, typically implemented in a tiled fashion, processing blocks of the output matrix in parallel. Consequently, reducing the linear layers into block diagonals does in principle not hinder GPU utilization, as the same tile-GEMMs can be used for the full product (1) or the block-diagonal (2) (with the latter requiring half as many). We use the PyTorch (Paszke et al., 2019) compiler to obtain efficiently running networks, but writing custom GPU-code would likely further improve the throughput.
(a) DeiT III.
(b) ConvNeXt-ISO and ResMLP.
Figure 5: Validation accuracy on ImageNet-1K versus model complexity as measured by the number of FLOPs required per image for a forward pass. Letters indicate the respective model sizes, see Tables 1 and 3 for details.
@@ 6.1 Results and Discussion
In prior work, image classification is often contextualized as a trade-off between accuracy and measures of model complexity, such as FLOPs and number of parameters. To that end, in Table 1, we highlight both the theoretical complexity, i.e. model size, and FLOPs, as well as the empirical speed as measured by the throughput (images processed per second), and the Top-1 classification accuracy on ImageNet-1K. Our results demonstrate that as we increase the scale of the networks, flopping-equivariant networks achieve competitive classification accuracy while requiring half the FLOPs, as highlighted in Figure 5, and showcase higher throughput compared to baseline implementations. The hybrid model $(\text{ViT-H})$-H ) achieves the best accuracy of all models, aligning with the results for hybrid models by Weiler & Cesa (2019). One reason the smaller equivariant models underperform in terms of accuracy is that they are quite limited in the number of parameters—the regularization methods in the training recipes for the corresponding non-equivariant models (with more parameters) may be too heavy. We leave the compute-heavy task of optimizing the training recipes for future work, as we believe that our experiments clearly demonstrate the main point of this paper—that equivariant networks can be scalable.
As shown in Figure 6, the relative throughput improvement of the equivariant networks becomes more pronounced as the size of the model increases, which can be attributed to a higher proportion of time spent on the linear layers, with the computational overhead from less optimized kernels becoming less significant. This is best illustrated by the superior speed of the equivariant ResMLP models for which linear layers dominate the computation. Our networks have no throughput improvement for attention or depthwise convolutions, which, despite not contributing the most FLOPs, are sometimes comparable in runtime to the dense layers. Future more efficient implementations of these layers, and further scaling of the embedding dimension, would pronounce the gain in throughput for the equivariant models.555 The reader will notice a flattening out of the rate of improvement between ViT-L and -H. ViT-H uses patch size $14\times 14$ while the other ViTs use patch size $16\times 16$. Smaller patch sizes means more patches and more compute in the attention layers.
6.2 Experimental Limitations
[…]
Table 1: Classification with Imagenet1k training. We contrast our flopping-equivariant networks to the originals using the same training recipes. The equivariant version of architecture X is denoted $\mathcal{E}(\text{X})$. $\mathcal{H}(\text{X})$ is a hybrid model with equivariant layers for the first half of the network. The baselines are not rerun, instead we show the results reported in their respective papers. The throughput and peak memory are measured on a single A100-40GB GPU with batch size fixed to 64 and compiled networks running mixed precision forward passes with no gradients.
@@ 6.2 Experimental Limitations
MLP-Mixers (ResMLP (Touvron et al., 2023))
@@
Vision Transformers (DeiT III (Touvron et al., 2022))
@@
Convolutional Networks (ConvNeXt (Liu et al., 2022))
@@ 7 Conclusion
The considered task of upright image classification is in a sense the least interesting for equivariant networks as the symmetry group is small and the output is invariant. We believe that future efforts should be directed towards equivariant networks as general visual feature extractors. Classic and recent works have already demonstrated the utility of having equivariant features for downstream vision tasks (Lowe, 2004; Loy & Eklundh, 2006; Lee et al., 2023; Bökman et al., 2024; Garrido et al., 2024).
@@ Appendix A Groups, Representations and Schur’s Lemma
This section is not required reading to understand the main takeaway of the paper. It serves to describe how the content of Section 3 can be formalized and generalized using group theory. The theory content of this section is common knowledge in geometric deep learning (Bronstein et al., 2021; Gerken et al., 2023; Weiler et al., 2023) and part of standard representation theory courses (Serre, 1977). Our contribution is to highlight the computational aspect more than has perhaps been done in previous work.
@@
Following the geometric deep learning blueprint (Bronstein et al., 2021), to design flopping-invariant neural networks $f=f_{N}\circ\cdots\circ f_{2}\circ f_{1}$, we enforce each network layer $f_{i}$ to be flopping-equivariant. The last layer is designed be flopping invariant, which is the special case of (11) where $\rho_{2}(g)\equiv I$. Designing an invariant neural network hence involves a choice of intermediate group representation $\rho_{i}$ for the output of each layer. We can write down all possible choices of $\rho_{i}$ in terms of fundamental building blocks called irreducible representations.
@@
Let $\varrho_{1}$ and $\varrho_{2}$ be two real irreps of a group $G$.
@@
The dimension in case 2 depends on whether the irrep $\varrho_{1}$ is of so-called real, complex or quaternary type (Serre, 1977).
@@
DeiT III (Touvron et al., 2022) ResMLP (Touvron et al., 2023) ConvNeXt (Liu et al., 2022)
@@ 3 Augment
MLP-Mixers (ResMLP (Touvron et al., 2023))
@@
Vision Transformers (DeiT III (Touvron et al., 2022))
@@
Convolutional Networks (ConvNeXt (Liu et al., 2022))
@@ Appendix B Experimental Setting
Software versioning. Our experiments build upon PyTorch (Paszke et al., 2019) and the timm (Wightman, 2019) library. We enable mixed-precision training using the deprecated NVIDIA library Apex, this is to mirror the training recipes of the benchmarks as closely as possible. To enable PyTorch’s compiler, we use a modern version ($\geq 2.0$). Specifically, we use PyTorch 2.5.1 with CUDA 11.8.
@@
    Flash Attention (Dao, 2023), used by the ViTs.
@@ Appendix C ResMLP Linear Layers Over Patches
This calculation is an example of a Clebsch-Gordan product or tensor product of representations, which has been used in some prior designs of equivariant networks (Kondor et al., 2018).